A psychologist wanted to test the belief that eating fish makes one smarter. she took a random sample of 12 persons took them through a fish oil supplement for one year and then assessed their IQ. her results were as follows 116,111,101,120,99,94,106,115,107,101,110,92. test at 5% the hypothesis that the fish oil supplement made a significant increase in the average IQ. The average IQ is expected to be 100.


1
Expert's answer
2021-07-03T09:26:32-0400
"mean =\bar{x}=\dfrac{1}{12}(116+111+101+120+99"

"+94+106+115+107+101+110+92)=106"

"Variance=s^2=\dfrac{1}{12-1}((116-106)^2"

"+(111-106)^2+(101-106)^2+(120-106)^2"

"+(99-106)^2+(94-106)^2+(106-106)^2"

"+(115-106)^2+(107-106)^2+(101-106)^2"

"+(110-106)^2+(92-106)^2)=78"

"s=\sqrt{s^2}=\sqrt{78}\approx8.83176"

The following null and alternative hypotheses need to be tested:

"H_0:\mu\leq100"

"H_1:\mu>100"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.05,"

"df=n-1=12-1=11" degrees of freedom, and the critical value for a right-tailed test is "t_c=1.795885."

The rejection region for this right-tailed test is"R=\{t:t>1.795885\}."

The t-statistic is computed as follows:

"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{106-100}{8.83176/\sqrt{11}}"

"\approx2.353394"

Since it is observed that "t=2.353394>1.795885=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value for "\alpha=0.05, df=11," "t=2.353394," right-tailed is "p=0.019129," and since "p=0.019129<0.05=\alpha," t is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than "100," at the "\alpha=0.05" significance level.